OpenStudy (anonymous):

What is the value of the expression? 14 + 16 ÷8 ---------- 2

OpenStudy (anonymous):

\[\frac{ 14+16\div8 }{ 2 } \] first \(16 \div 8=2\) \[\frac{ 14+\color{red}2 }{2 }\]

OpenStudy (anonymous):

\[\frac{ 16 }{ 2 }=8\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

the order of operations is important, first brackets, then factors (roots count as factor too), then multiplication (division counts as multiplication too), then adding (subtracting counts as adding too)

OpenStudy (anonymous):

in a fraction, brackets are implied, by the way, that's why 14+2 goes before dividing, even though division is before adding normally